Charge is placed on a metallic sphere which is surrounded by air. If the radius of the sphere is 15 cm, how much charge can be placed on the sphere before the air near the sphere suffers electric breakdown? The critical electric field strength that leads to breakdown in air is 3 x 10^6 N/C


1
Expert's answer
2022-02-28T10:05:57-0500

The electric field

"E=k\frac{q}{r^2}"

Hence, the required charge

"q=Er^2/k\\
=3*10^6*0.15^2/9*10^9=7.5*10^{-6}\:\rm C"